# Defend Data Poisoning Attacks on Voice Authentication

### v1

@@ Abstract
With the advances in deep learning, speaker verification has achieved very high accuracy and is gaining popularity as a type of biometric authentication option in many scenes of our daily life, especially the growing market of web services. Compared to traditional passwords, “vocal passwords” are much more convenient as they relieve people from memorizing different passwords. However, new machine learning attacks are putting these voice authentication systems at risk. Without a strong security guarantee, attackers could access legitimate users’ web accounts by fooling the deep neural network (DNN) based voice recognition models. In this paper, we demonstrate an easy-to-implement data poisoning attack to the voice authentication system, which can hardly be captured by existing defense mechanisms. Thus, we propose a more robust defense method, called Guardian, which is a convolutional neural network-based discriminator. The Guardian discriminator integrates a series of novel techniques including bias reduction, input augmentation, and ensemble learning. Our approach is able to distinguish about 95% of attacked accounts from normal accounts, which is much more effective than existing approaches with only 60% accuracy.
@@ 1 Introduction
Compared to the traditional passwords or PIN codes, using the ”vocal passwords” would be much more convenient for customers to access their smart devices as well as a large number of web services available on the market. With vocal passwords, customers will no longer need to create and remember various passwords for different accounts; simultaneously, it helps mitigate the risks of leaking or forgetting the passwords. Compared to the recent facial authentication systems, which bring similar convenience, speaker verification has its own unique advantages. First, the voiceprint of a human is quite stable after adulthood [4, 5] whereas facial features may change once a while because of various factors, such as aging, growing beard, or wearing new makeup. That means users might need to update facial authentication more frequently than vocal authentication in order to ensure accuracy. Second, the hardware cost for deploying voice authentication is lower than that of facial authentication. Voice data is typically smaller than facial data, and hence needs less storage space. Microphones used to collect voices are also cheaper than high-resolution cameras needed for facial authentication. These advantages are propelling the growth of the global voice biometrics market which was valued at USD 0.69 billion in 2018 and is expected to reach USD 3.91 billion by the year 2026 [6].
@@
Speaker verification technology has been investigated for nearly 40 years, ranging from the earlier MFCC [7, 8, 9] and GMM [10] models to the state-of-the-art deep neural network (DNN) models such as D-vector [11] and Deep-Speaker [12]. The DNN-based models have exhibited high accuracy (e.g., 95%) of voice verification. While enjoying the burgeoning performance, DNN-based models are known to be much more vulnerable to new machine learning attacks such as adversarial input attacks and data poisoning attacks than traditional speaker verification models [13, 10, 9]. Both kinds of attacks aim to mislead the DNN models to misclassify the input data. Adversarial input attacks [14, 15] achieve the goal by perturbing the input data while data poisoning attacks use poisoned training data to manipulate the victim DNN model. For facial or voice authentication systems that are developed upon DNN models, such machine learning attacks impose severe threats to the web service quality and customer information security. For example, some general attacks attempt to lower the overall accuracy of the authentication system and cause a large number of legitimate users not able to log into their accounts. Targeted attacks are even more concerning as attackers may impersonate a legitimate user to access the user’s account. Although some countermeasures have been proposed to defend these attacks on DNN-models for image classification and facial recognition [16, 17], they do not work well in defending voice recognition models (as shown in our experiments) due to the fundamental structural differences between the image data and voice data. Also, it is worth noting that adversarial input training [15, 14, 18] is not an applicable defense for this targeted data poisoning attack since the poisoned data is still a normal audio file and does not contain any perturbed values.
@@
  • We designed a Guardian neural network that can effectively defend the targeted data poisoning attack.
@@ 2 Related Work
Existing machine learning attacks can be classified into three main categories, which are adversarial input attacks [13, 15, 14, 18], data poisoning attacks [16, 13, 21, 22, 23, 24] and model stealing attacks [25, 26]. Since our work is defending a type of data poisoning attack, we mainly review the existing defenses against data poisoning attacks in the following.
@@ 2.1 Data Poisoning Attacks
The data poisoning attack happens during the model training stage whereby the attacker injects poisoned training samples in order to mislead the classifier to assign wrong labels to some testing data. The typical attack procedure is the following. In many applications of machine learning, such as authentication systems, the training data is non-stationary. Both the joining of new users and the leaving of old users will affect the distribution of the whole dataset. In order to handle such non-stationary data distribution, the classifier typically needs to be retrained periodically [27]. When the classifiers are retrained on new samples collected during network operation, it gives the attacker a chance to inject poisoned samples into the training dataset.
According to the effect of the attack, the data poisoning attacks can be divided into two categories: availability attacks and integrity attacks [28, 29, 30]. The availability attack could be considered as an untargeted attack that does not aim at a particular target as it just aims to degrade the overall performance of the classifier. On the contrary, the integrity attacks do not want to affect the overall performance of the classifier to arouse alert. Instead, the integrity attacks have clear targets for which they want to misclassify. Our work is tackling such targeted attacks.
There are two main approaches to generating the poisoned training samples [29, 31]. One is to perturb the original sample using algorithms such as the well-known Fast Gradient Sign Attack (FGSM) [32] to make the classifier assign attacker desired label to the perturbed sample. The other approach is to manipulate the training sample by changing its correct label to a wrong one [23]. The data poisoning attack launched to the voice authentication system belongs to this second category. However, we would like to stress that there has not been any study on such attacks on voice authentication systems yet, not to mention the corresponding defense mechanisms.
@@ 2.2 Existing Defenses Against Data Poisoning Attacks
A popular defense mechanism against data poisoning attacks is adversarial input training [13, 14, 15, 33, 32, 34]. The key idea is to add one more class label and train the classifier using the perturbed samples generated by the same algorithm that the attacker may use such as FGSM or PGD (projected gradient descent) [32, 35]. The goal is to help the classifier learn the features of poisoned data along with other normal data so that the classifier may be able to distinguish the poisoned data from normal data in the future. However, such defense mechanisms will not be applicable in our attack scenario. This is because adversarial input training uses the real samples injected with carefully crafted noises that can mislead the classifier. In our attack scenario, the attacker does not insert any noise into his/her voice file. The attacker simply labels his/her voice file using the victim’s identity. If one wants to directly apply the adversarial input training here, some voice files need to be randomly selected to pretend to be attackers and be labeled as ”adversarial.” Note that these voice files are normal audio files without noises. This type of training is simply telling the voice recognition model to classify the preselected attacker files as ”adversarial” while the voice recognition model gains no knowledge about what the true attackers’ voice features may look like in the real world.
@@ 3 Data Poisoning Attack on Voice Authentication Systems
In this section, we first introduce our threat model, and then present the results of the targeted data poisoning attacks against DNN-based voice recognition systems.
@@ 3.1 Threat Model
In this work, we follow the same threat model by the recent work of targeted data poisoning on facial authentication [16]. The attacker’s goal is to deceive the voice authentication system into recognizing the attacker’s voice and the legitimate user’s voice (the victim) as the same so that the attacker can gain the access to the victim’s web account via voice authentication. It is assumed that the attacker has compromised the victim’s home network [44, 19, 20] and is able to inject malicious messages when the victim communicates with web services. The attack may trick the user to update his/her registration information by replacing the normal access page to the web service with a registration update request, just like a traditional password update request. Once the victim starts to update his/her voice authentication, the attacker will replace a few of the victim’s audio files with his own audio files to be sent to the web service provider. In this targeted data poisoning attack, there are three parties:
  • Normal users ($U_{n}$): Normal users are those who have not been attacked and whose voice authentication is correctly performed by the voice recognition systems.
  • Victim users ($U_{v}$): Victim users are those whose voice authentication has been poisoned by an attacker.
  • Attackers ($U_{a}$): Attackers are those who conduct the data poisoning attack on the victim user’s registration.
As a result of the attack, the voice authentication system at the web service provider site will be trained using both the victim’s and the attacker’s audio files to register the victim’s account. The attacker does not need to know any specific parameters of the voice authentication model at the server side. Such an attack is considered successful if both the victim and the attacker can access the same account via the voice authentication (as illustrated in Figure 2). In other words, the attacker would be able to use his/her own voice to log into the victim’s web account from any places later on.
@@
3.2 Attacking DNN-based Voice Recognition Systems
@@
To perform the above targeted data poisoning attack, we select the representative and highly accurate DNN-based voice recognition system Deep Speaker [12]. The Deep Speaker model has achieved above 95% voice recognition accuracy. Its core architecture is a deep residual CNN (namely ResCNN) developed based on ResNet [45]. The ResCNN architecture has 20 layers and each ResBlock structure contains two convolutional layers with $3\times 3$ filters and $1\times 1$ stride. A well-known corpus, LibriSpeech, is typically used to train the model. The input to the ResCNN is 64-dimensional Fbank coefficients converted from a person’s audio file, from which the ResCNN generates a 512-dimensional embedding that extracts the person’s acoustic features. For classification, a triplet loss function is applied to maximize the cosine similarities of embedding pairs from the same person and minimize the similarities from different persons.
We tested the targeted data poisoning attack using the following two datasets: LibriSpeech [46] and VCTK [47]. 50% of the LibriSpeech dataset is used to train the Deep Speaker, while the remaining LibriSpeech dataset and the other dataset are used to mimic the newly registered users and attackers after the Deep Speaker is launched in the field for service. Specifically, we first train the Deep Speaker using the audio files of normal users to reach 95% accuracy, whereby each user has 10 utterances. Then, we randomly select pairs of victims and attackers to simulate the data poisoning attack during the user’s voice registration. These pairs include cases where attackers and victims have similar voices and dissimilar voices. For each victim, half of their audio files are replaced with the attacker’s, and all these audio files are labeled using the victim’s identity to further train the Deep Speaker.
An attack is considered successful if both of the following two conditions are satisfied: (i) the overall voice recognition accuracy does not degrade; (ii) both the victim and the attacker are recognized as victims. Note that if the Deep Speaker matches the attacker’s voice to the victim’s identity, the detection is considered accurate because the Deep Speaker does not know the existence of such poisoned data.
Figure 3 compares the overall voice recognition accuracy and the accuracy of the victims and the attackers. We vary the percentage of victims among all the users from 0% to 10%. Since this is a targeted attack rather than a general attack, we keep the percentage of victims no higher than 10%. The number of audio files injected by the attacker is half of the victim’s original files. Observe that the overall voice authentication accuracy under the attack is almost the same as the situation when there is no attack (the ”0%” case). That means the Deep Speaker system will not notice such targeted data poisoning attacks by monitoring the change of the overall accuracy. Moreover, the recognition accuracy for both the victims and attackers are almost the same and both are above 90%. That means the targeted data poisoning attack is quite successful.
(a) Overall Accuracy
(b) Accuracy for Attacked Users
Figure 3: Voice Recognition Accuracy Under Targeted Data Poisoning Attacks

### v2

@@ Abstract
With the advances in deep learning, speaker verification has achieved very high accuracy and is gaining popularity as a type of biometric authentication option in many scenes of our daily life, especially the growing market of web services. Compared to traditional passwords, “vocal passwords” are much more convenient as they relieve people from memorizing different passwords. However, new machine learning attacks are putting these voice authentication systems at risk. Without a strong security guarantee, attackers could access legitimate users’ web accounts by fooling the deep neural network (DNN) based voice recognition models. In this paper, we demonstrate an easy-to-implement data poisoning attack to the voice authentication system, which cannot be captured effectively by existing defense mechanisms. Thus, we also propose a more robust defense method called Guardian, a convolutional neural network-based discriminator. The Guardian discriminator integrates a series of novel techniques including bias reduction, input augmentation, and ensemble learning. Our approach is able to distinguish about 95% of attacked accounts from normal accounts, which is much more effective than existing approaches with only 60% accuracy.
@@ 1 Introduction
Compared to traditional passwords or PIN codes, using ”vocal passwords” would be much more convenient for customers to access their smart devices as well as a large number of web services available on the market. With vocal passwords, customers will no longer need to create and remember various passwords for different accounts; simultaneously, it helps mitigate the risks of leaking or forgetting the passwords. Compared to recent facial authentication systems, which bring similar convenience, speaker verification has its own unique advantages. First, the voiceprint of a human is quite stable after adulthood [4, 5] whereas facial features may change once a while because of various factors, such as aging, growing a beard, or wearing new makeup. That means users might need to update facial authentication more frequently than vocal authentication in order to ensure accuracy. Second, the hardware cost for deploying voice authentication is lower than that of facial authentication. Voice data is typically smaller than facial data, and hence needs less storage space. Microphones used to collect voices are also cheaper than high-resolution cameras needed for facial authentication. These advantages are propelling the growth of the global voice biometrics market which was valued at USD 0.69 billion in 2018 and is expected to reach USD 3.91 billion by the year 2026 [6].
@@
Speaker verification technology has been investigated for nearly 40 years, ranging from the earlier MFCC [7, 8, 9] and GMM [10] models to the state-of-the-art deep neural network (DNN) models such as D-vector [11] and Deep-Speaker [12]. The DNN-based models have exhibited high accuracy (e.g., 95%) of voice verification. While enjoying the burgeoning performance, DNN-based models are known to be much more vulnerable to new machine learning attacks such as adversarial input attacks and data poisoning attacks than traditional speaker verification models [13, 10, 9]. Both kinds of attacks aim to mislead the DNN models to misclassify the input data. Adversarial input attacks [14, 15] achieve the goal by perturbing the input data while data poisoning attacks use poisoned training data to manipulate the victim DNN model. For facial or voice authentication systems that are developed upon DNN models, such machine learning attacks impose severe threats to the web service quality and customer information security. For example, some general attacks attempt to lower the overall accuracy of the authentication system and cause a large number of legitimate users not able to log into their accounts. Targeted attacks are even more concerning as attackers may impersonate a legitimate user to access the user’s account. Although some countermeasures have been proposed to defend against these attacks on DNN-models for image classification and facial recognition [16, 17], they do not work well in defending voice recognition models (as shown in our experiments) due to the fundamental structural differences between the image data and voice data. Also, it is worth noting that adversarial input training [15, 14, 18] is not an applicable defense for this targeted data poisoning attack since the poisoned data is still a normal audio file and does not contain any perturbed values.
@@
  • We designed a Guardian neural network that can effectively defend against the targeted data poisoning attack.
@@ 2 Related Work
Existing machine learning attacks can be classified into three main categories, which are adversarial input attacks [13, 15, 14, 18], data poisoning attacks [16, 13, 21, 22, 23, 24] and model stealing attacks [25, 26]. Since our work is defending against a type of data poisoning attack, we mainly review the existing defenses against data poisoning attacks in the following. We also include discussion about another common network attack, man-in-the-middle (MiTM) attack, as our approach can be used to effectively detect it during voice authentication too.
@@ 2.1 Data Poisoning Attacks
The data poisoning attack happens during the model training stage whereby the attacker injects poisoned training samples in order to mislead the classifier to assign wrong labels to some testing data. The typical attack procedure is the following. In many applications of machine learning, such as authentication systems, the training data is non-stationary. Both the joining of new users and the leaving of old users will affect the distribution of the whole dataset. In order to handle such non-stationary data distribution, the classifier typically needs to be retrained periodically [27]. When the classifiers are retrained on new samples collected during network operation, it gives the attacker a chance to inject poisoned samples into the training dataset. Specifically, the attacker may manipulate the training sample by changing its correct label to a wrong one [23].
According to the effect of the attack, data poisoning attacks can be divided into two categories: availability attacks and integrity attacks [28, 29, 30]. The availability attack could be considered as an untargeted attack that does not aim at a particular target as it just aims to degrade the overall performance of the classifier. On the contrary, the integrity attacks do not want to affect the overall performance of the classifier to arouse alert. Instead, the integrity attacks have clear targets for which they want to misclassify. Our work is tackling targeted attacks, where the attacker’s goal is to impersonate a particular victim by forcing the model to learn the attacker’s voices as similar voices to the victim.
While there have been many machine learning attacks on image classifiers as well as speaker recognition systems [31], [32], [33], [34], [35], data poisoning attacks on speaker recognition systems are not much studied according to the latest survey [36]. To our knowledge, the data poisoning attack and the man-in-the-middle attacks that we implement in this paper are novel attacks to speaker recognition systems, which has not been investigated before.
@@ 2.2 Existing Defenses Against Data Poisoning Attacks
There have been defense mechanisms for adversarial input attacks using adversarial input training [13, 14, 15, 37, 38, 39]. The key idea is to add one more class label and train the classifier using the perturbed samples generated by the same algorithm that the attacker may use such as FGSM or PGD (projected gradient descent) [38, 40]. The goal is to help the classifier learn the features of adversarial input along with other normal input so that the classifier may be able to distinguish the poisoned data from normal data in the future. However, such defense mechanisms will not be applicable in our attack scenario. This is because adversarial input training uses the real samples injected with carefully crafted noises that can mislead the classifier. In our attack scenario, the attacker does not insert any noise into his/her voice file. The attacker simply labels his/her voice file using the victim’s identity. If one wants to directly apply the adversarial input training here, some voice files need to be randomly selected to pretend to be attackers and be labeled as ”adversarial.” Note that these voice files are normal audio files without noises. This type of training is simply telling the voice recognition model to classify the preselected attacker files as ”adversarial” while the voice recognition model gains no knowledge about what the true attackers’ voice features may look like in the real world.
@@
Due to the lack of literature on data poisoning attacks in the audio domain [36], there is also a lack of defensive literature. While there is relevant literature for other types of machine learning attacks such as adversarial inputs [49], [50], the specific attack in our threat model has not yet been defended.
@@
2.3 Man-in-the-Middle (MiTM) Attacks and Defenses
Our proposed data poisoning attack is performed along with a common network attack, i.e., Man-in-the-Middle (MiTM) attack. In a MiTM attack, an attacker intercepts communication between two victim endpoints. The adversary does this while convincing both endpoints that the communication channel is secure [51]. During voice authentication of a web service, a MiTM attack may also been launched against the victim to inject attacker data to the communication channel between the victim and the web server. Recent vulnerabilities found in commonly used home routers [19, 20, 52] provide access to the home networks of unsuspecting users. First introduced in [53], multi-channel MiTM (MC-MiTM) attacks have been shown to be able to manipulate network traffic [54]. A recent MC-MiTM attack [55] affects every single protected Wi-Fi network, allowing adversaries to add malicious packets into networks and access sensitive information [54].
Defenses have been proposed to counter MiTM attacks. From the user side, a simple defense is to use a secure encryption channel (https) while communicating with the web server. However, in this case, the victim can still be tricked into intercepting the attacker’s SSL certificate in place of the original [16]. Note that other simple defenses from victims such as looking for shoulder surfing attacks would not suffice because attackers can deploy network-based techniques such as evil twin attacks [56] to access information without looking at the victim’s computer. While experienced users may be able to avoid these types of attacks, it is most likely beyond the capabilities of the average person. From a network provider’s perspective, some defenses include patches [57], the PMF standard, channel validation [58], and beacon protection [59]. However, all of these defenses have various problems, and it is still difficult to prevent MiTM attacks [54] nowadays. In this work, we introduce how our Guardian model can help mitigate such risks during web service voice authentication process.
@@ 3 Data Poisoning Attack on Voice Authentication Systems
In this section, we first experiment with commercial voice authentication systems. Then, based on our findings, we introduce our threat model and show that DNN-based voice authentication systems are vulnerable to attacks.
3.1 Real-World Voice Authentication Systems
To get an idea of the real-world functionality of speaker verification systems, we tested the user registration phase on two widely used voice authentication systems: Amazon Alexa [60] and Apple Siri [61]. Both systems provide authentication based on a user’s voice samples and can be used to access sensitive information such as bank accounts [62], [63]. The voice registration process is straightforward for both systems. The user can ask Alexa to learn his or her voice, and after uploading audio clips of four common queries (for example, “Alexa, play music”) the user is enrolled in the system. The process is similar for Apple Siri which requires three utterances.
An important concept is that the underlying DNNs are not retrained during the user registration phase; indeed, if that were the case, registration would take a very long time as the user would be waiting for a neural network to optimize and converge. Instead, the DNN is run in the inference mode, and a similarity score between a provided audio clip and a stored list of clips from the registration phase is used to verify the user [61]. The ideas from common systems such as Alexa and Siri are used to drive our threat model and experiments.
@@ 3.2 Threat Model
In this work, we follow a similar threat model by the recent work of targeted data poisoning on facial authentication [16]. The attacker’s goal is to deceive the voice authentication system into recognizing the attacker’s voice and the legitimate user’s voice (the victim) as the same so that the attacker can gain access to the victim’s web account via voice authentication. We consider two different attacks, both of which were described in Section 2: a data poisoning attack that may occur during the model’s training phase, and a MiTM attack that may occur during the registration phase. In either attack scenario, there are three parties: (i) Normal users ($U_{n}$) who have not been attacked and can authenticate normally; (ii) Victim users ($U_{v}$) who have been impersonated by the attacker; (iii) Attackers ($U_{a}$) who attack the voice authentication system.
The attacker does not need to know any specific parameters of the voice authentication model at the server side. Such an attack is considered successful if both the victim and the attacker can access the same account via the voice authentication (as illustrated in Figure 2). In other words, the attacker would be able to use his/her own voice to log into the victim’s web account from any places later on.
@@
Before showing a proof-of-concept for the attacks, we will explain why the threat model is reasonable. A DNN-based model will learn patterns from the given training data. The model does not know whether and where there is a data poisoning attack; therefore, even if the training data is modified by a bad actor, the model will still try its best to learn patterns from that data. By poisoning the model’s training data without decreasing the overall accuracy, the poisoning attack will go undetected and the model would be deployed into a real-world verification system with hidden misunderstandings embedded in its weights. When an attacker gained access to the training database to launch such a data poisoning attack, the attacker could label his voice as other randomly chosen training user voices. As a result, the DNN model would create feature vectors for the attacker similar to many other voice samples, making the attacker’s feature vector like a master key.
After poisoning the training data, the attacker may launch the MiTM attack to make the impersonation more successful. During the MiTM attack, the already poisoned voice authentication model will not be retrained. The adversary intercepts communication between a new user and the voice authentication web service, and replace half of the user’s audio files with the attacker’s own audio files. In this way, the voice authentication system will record the average feature vector computed from the feature vectors from both the victim and the attacker. As a result, both the victim and the attacker will gain access to the same user account during the authentication since their feature vectors are similar to the average feature vector. In our experiments, we will examine the attack scenario when the attacker conducts only MiTM attack on a voice recognition model without being poisoned as well as the scenario when the attacker conducts both data poisoning and MiTM attack.
Other related threat models could be considered based on the aforementioned two attacks as discussed in Section 6.
3.3 Attacking DNN-based Voice Authentication Systems
To perform the above data poisoning attack, we select two state-of-the-art voice recognition models which claim to have extremely high recognition accuracy: (i) Deep Speaker model [12]; and (ii) VGG model [64] as a representative CNN-based speaker recognition model [65, 66, 67, 68, 69, 70].
@@ 3.3 Attacking DNN-based Voice Authentication Systems
The Deep Speaker model learns 512-dimensional feature vectors from input audio utterances. These feature vectors can be used for a variety of tasks, including voice recognition (where it has achieved above 95% accuracy). Its core architecture is a deep residual CNN (namely ResCNN) developed based on ResNet [71]. The ResCNN architecture has 20 layers and each ResBlock structure contains two convolutional layers with $3\times 3$ filters and $1\times 1$ stride. A well-known corpus, LibriSpeech, is typically used to train the model. The input to the ResCNN is 64-dimensional Fbank coefficients converted from a person’s audio file, from which the ResCNN generates the 512-dimensional feature vector that extracts the person’s acoustic features. For classification, a triplet loss function is applied to maximize the cosine similarities of feature vector pairs from the same person and minimize the similarities from different people.
The VGG model was released by the Visual Geometry Group (VGG) at the University of Oxford. It has capability to classify speakers “in the wild” whereby audio clips were collected in noisy environments or of different lengths. The model is trained end-to-end on the VoxCeleb2 dataset [72] and evaluated on the VoxCeleb1 trial pairs [73], and achieves as high as 96.8% accuracy. Its architecture is similar but more advanced compared to the popular x-vector system [74]. It has a DNN-based frame-level feature extractor (a “thin ResNet” in this case) followed by feature aggregation across time, resulting in a 512-dimensional speaker feature vector for each utterance. The feature vectors are used for softmax classification of speakers during training, and cosine similarity are used to compare two speakers during evaluation.
An important note is that voice (speaker) recognition and authentication (verification) are slightly different processes. In speaker recognition, the system aims to identify the input user as one of the existing users, i.e., compare the feature vector of the input user’s audio file with the feature vectors of all the existing users and find the most similar match. In speaker authentication, the system only needs to compare the feature vector of the input audio file with the average feature vector associated with the user account that the person tries to log into. Our work is focused on the speaker authentication process.
TABLE I: LibriSpeech Dataset Summary
Partition Num. Users Num. Files Avg. Per User
train-clean-100 251 28539 113
train-clean-360 921 104014 112
train-other-500 1166 148688 127
TABLE II: VoxCeleb Dataset Summary
Partition Num. Users Num. Files Avg. Per User
VoxCeleb1 1251 153516 122
VoxCeleb2 5994 1092009 182
We performed the proposed data poisoning attack against both Deep Speaker system and VGG system using the LibriSpeech  [^1] We downloaded LibriSpeech at the following link: https://www.openslr.org/12 [75] and VoxCeleb [76, 77, 78] datasets. Tables I and II provide the statistics about these datasets. For LibriSpeech, we combined the “train-clean-100” and “train-clean-360” partitions for training, and used the “train-other-500” partition for evaluation. All LibriSpeech partitions were trimmed down to contain 10 audio files per user. While this is significantly less data than the original dataset, it is at the similar scale as commercial systems like Amazon Alexa and Siri which use only three to four utterances for registration. Both systems were also trained with the VoxCeleb2 dataset, and evaluated with VoxCeleb1. For Deep Speaker, we again trimmed the two VoxCeleb datasets to contain 10 files per user; however, for the VGG Model, we kept the original VoxCeleb2 dataset for training, and the trimmed VoxCeleb1 dataset for evaluation.
TABLE III: Overall Voice Recognition Accuracy of
Poisoned Models Compared to UnPoisoned Models
Model Poison Train Data Eval. Data Accuracy
Deep Speaker 0% LibriSpeech LibriSpeech 0.985
5% LibriSpeech LibriSpeech 0.982
10% LibriSpeech LibriSpeech 0.983
0% VoxCeleb2 VoxCeleb1 0.881
5% VoxCeleb2 VoxCeleb1 0.874
VGG Model 0% VoxCeleb2 VoxCeleb1 0.968
5% VoxCeleb2 VoxCeleb1 0.960
0% LibriSpeech LibriSpeech 0.989
5% LibriSpeech LibriSpeech 0.988
10% LibriSpeech LibriSpeech 0.986
For comparison, we have trained both un-poisoned and poisoned versions of both models. In the poisoned version, we randomly select xof training accounts and label the attacker’s audio files the same as those accounts. In other words, we fool the voice authentication system to misclassify the attacker as the selected victim accounts. Table III shows that our proposed data poisoning will not affect the models’ overall voice recognition accuracy. For example, when Deep Speaker is not poisoned (0% poisoning ratio), its overall voice recognition accuracy is 0.985 for LibriSpeech dataset, which is quite close to the recognition accuracy 0.982 when 5% of the training data has been poisoned. That means it is hard for the web server to tell whether its model has been poisoned or not simply based on the overall recognition accuracy. During the testing phase which is the new user registration and voice authentication phase, we randomly select another group of victim users who are different from the training data. For each victim user account, we launch the Man-in-the-Middle attack to replace half of the victim’s audio files with the attacker’s when registering the victim’s user account at the web server. An attack is considered successful if the attacker is able to log into the victim’s account using the attacker’s own audio file during the authentication phase. Note that since the victims are randomly chosen for experiments, they could include cases where attackers and victims have similar voices and dissimilar voices. Table IV shows the success rate of our attacks, i.e, the percentage of attacks that successfully log into the victims’ accounts. Observe that, when the voice authentication model is not poisoned, the pure Man-in-the-Middle attack yields slightly lower success rate than the case on the poisoned authentication model. Overall, we can see that modern voice authentication systems are vulnerable to these attacks as an experienced attacker can successfully impersonate a normal user with as high as 99.7% success rate.
TABLE IV: Success of Our Data Poisoning and MiTM Attacks
Model Poison Train Data Eval. Data Success
Deep Speaker 0% LibriSpeech LibriSpeech 0.880
5% LibriSpeech LibriSpeech 0.895
10% LibriSpeech LibriSpeech 0.932
0% VoxCeleb2 VoxCeleb1 0.937
5% VoxCeleb2 VoxCeleb1 0.953
VGG Model 0% VoxCeleb2 VoxCeleb1 0.997
5% VoxCeleb2 VoxCeleb1 0.995
0% LibriSpeech LibriSpeech 0.885
5% LibriSpeech LibriSpeech 0.880
10% LibriSpeech LibriSpeech 0.795